# More Credible, Comprehensive and Challenging

### v1

@@ Abstract
[…]
@@
[…]
@@
Table 1: (a) Re-annotation: For the test or validation sets of four datasets, the proportion of updated examples ranges from 9.7% to 30%. (b) Re-evaluation: Silver denotes the Accuracy on the original dataset, while Gold represents the results on the re-annotated dataset. Rigorous revision and evaluation reveal LRMs’ actual performance and DeepSeek-R1’s significant improvement over DeepSeek-V3.
@@ 2.1 Updates to Public Datasets
Following prior work Krumdick et al. (2024); Zhao et al. (2024), we retain the format of questions with optional hybrid contexts as input, accompanied by Python-formatted solutions and program-executed numerical results. Due to the specialized nature of complex financial problems and the high cost of expert annotation, we observe certain limitations in existing datasets Chen et al. (2021b); Zhu et al. (2021); Zhao et al. (2024), including unsolvable problems, ambiguous questions (e.g., the phrasing “the range of” confuses the LLMs, as it is not clear whether to output a range like 70.18-81.05 or a specific difference of 10.87), oversimplified processes, incorrect answers, and relaxed evaluation criteria. Statistical details of these issues are presented in Table 1.
@@ 2.2 Function Library Construction
We begin by collecting 6,138 financial encyclopedia articles from Investopedia, a platform renowned for its extensive expertise in financial knowledge [^1] https://www.investopedia.com. Each article provides a detailed introduction to a specific financial term, covering fundamental concepts, application scenarios, and potential limitations, some including relevant calculation formulas and practical examples. To distill dense, structured financial reasoning knowledge while reducing annotation costs, we instruct GPT-4o to extract potential financial calculation functions from each article according to a predefined format. Each function is required to include a semantically meaningful signature, a concise and clear docstring (functionality, parameters, return values, applicable constraints, and other notes), and step-by-step implementation code with appropriate annotations. Finally, we organize financial experts to rigorously review and revise the generated functions, ensuring their professional expression and logical correctness. Detailed examples are illustrated in Figure 3 and Table 15.
@@ 2.3 Expansion of Data Annotation
Existing financial question-answering datasets (e.g., CodeFinQA, CodeTAT-QA) focus primarily on evaluating models’ basic concept understanding, precise numerical extraction, and simple calculation abilities within given contexts. Problem-solving processes in these datasets typically involve fewer reasoning steps (e.g., calculate the difference in net profit over two years). These datasets often suffer from redundancy in simple questions and a lack of complex questions, failing to adequately assess the reasoning capabilities of LRMs, such as knowledge application, constraint emphasis and long thought (e.g., compute the net return rate of a fund in Figure 1). As a result, the true reasoning capabilities of LRMs cannot be evaluated comprehensively and objectively. Therefore, optimizing data construction methods, rigorously verifying data quality, and building more challenging datasets have become crucial to improve the evaluation of financial reasoning tasks.
@@ 2.4 Data Quality Assurance
Table 2: Statistics of the function library and FinanceReasoning dataset (Avg values of three subsets).
@@ 2.5 Data Grading and Statistics
As illustrated in subsection 2.4 and Figure 4, based on the difficulty of reasoning, we divided the problems into three subsets: Easy (1,000 examples), Medium (1,000 examples), and Hard (238 examples). More analyses are in Appendix C.2. To promote the evaluation of LRMs’ reasoning capabilities in the financial domain, we have made all problems publicly available, along with the complete financial function library.
@@ 3 Evaluation System
We developed an evaluation system for complex reasoning on FinanceReasoning, where all evaluations of LLMs were conducted by calling their official API interfaces. LABEL:tab:model-detail illustrates exact model version we used.
@@ 3.1 Large Language Models
We focused on evaluating six of the most powerful large reasoning models currently available.
@@ 3.2 Evaluation Methods
Following Zhao et al. (2024), we evaluated LLMs with Chain-of-Thought Wei et al. (2022) and Program-of-Thought Chen et al. (2023b) to achieve optimal performance and make comparisons.
@@
We adopt the answer extraction pipeline from Zhao et al. (2024), using GPT-4o-mini to extract numerical results from the output under the CoT setting, and executing the program from the output under the PoT setting. Finally, we perform a strict accuracy evaluation comparing the numerical results with the ground truth within 0.2% error margin.
@@ 4.1 Main Results (RQ1)
The results demonstrate that the powerful LRM (i.e., OpenAI o1) using PoT prompting method achieves the best performance, with an accuracy of 89.1% on the Hard subset, significantly outperforming other LRMs and LLMs. On the Easy and Medium subsets, the evaluated LLMs achieve accuracy above 87%, except for the Llama models, where the advantage of LRMs is less pronounced. This further validates that simpler datasets have largely been solved by LLMs, making it difficult to assess the real reasoning capabilities of LRMs. On the Hard subset, LRMs exhibit a clear advantage over LLMs in CoT, further confirming the superiority of LRMs in complex reasoning tasks.
In particular, on the Hard subset, the current superior LRMs (i.e., OpenAI o1 and DeepSeek-R1) exhibit distinct performance contrasts in CoT and PoT settings. In the CoT setting, DeepSeek-R1 achieves a 2.1% higher accuracy than OpenAI o1 (83.2% vs. 81.1%). However, the PoT prompting method significantly enhances OpenAI o1’s performance, allowing it to surpass DeepSeek-R1 (89.1% vs. 85.3%). This suggests that DeepSeek-R1 excels at text-based step-by-step reasoning, while OpenAI o1 outperforms in programming capabilities. This discrepancy may be due to differences in their training methods and training data.
@@ 4.2 Error Analysis (RQ2)
To better analyze the capabilities and limitations of LRMs on difficult problems in our dataset, we conduct a thorough and comprehensive error analysis. This analysis is based on 80 DeepSeek-R1 failure cases under PoT, with stratified sampling (20 Easy /20 Medium /40 Hard). We summarize four types of error in the current LRMs on challenging domain-specific reasoning problems, some of which involve compound errors. The detailed error distribution are shown in Table 4. More details of error cases are provided in Appendix B.
@@
  • Formula Application Errors and Numerical Calculation Errors increase with difficulty, especially the former, which supports the effectiveness of our difficulty categorization. In contrast, Numerical Extraction Errors remain consistently low across all subsets. These findings suggest that the model’s domain-specific knowledge comprehension and application capabilities should be enhanced through retrieval-augmented generation (RAG) or reinforcement learning, thereby improving its performance on complex reasoning tasks in specialized domains.
@@ 4.3 Knowledge Augmentation (RQ3)
LLM-instructed Retrieval & LLM as Judge 91.60 (+8.41) 86.97 (+3.78)
Table 5: Results of different knowledge augmentation methods on the Hard subset of FinanceReasoning. GPT-4o with refined knowledge augmentation, outperforming OpenAI o1 (91.6% vs 89.1%) under PoT setting.
@@
As shown in subsection 4.3, the format and method of knowledge augmentation significantly affect the performance of the model reasoning. Specifically, LLMs enhanced with financial function knowledge significantly outperform those enhanced with passage knowledge, as financial functions serve as refined reasoning knowledge. Excessive and intricate passages can disrupt the model’s reasoning abilities, resulting in diminished performance for both LLMs and LRMs. Taking advantage of the improved retrieval efficiency caused by LLM-Instructed Knowledge Retrieval, the combination approach achieves the best performance, improving the accuracy of GPT-4o to 91.6% with PoT. More analyses are shown in Appendix E.
@@ 4.4 Reasoner with Programmer (RQ4)
Specifically, we employ DeepSeek-R1 as Reasoner, and the best-performing models under PoT settings, Claude 3.5 Sonnet and GPT-4o, as Programmers, respectively. As shown in Figure 5, the combination of models achieves significant improvements compared to individual models. Compared to GPT-4o, Claude 3.5 Sonnet demonstrates a stronger ability to follow the given reasoning logic and generate precise code without introducing noise, owing to its programming advantages. The combination of DeepSeek-R1 and Claude 3.5 Sonnet achieves an accuracy of 87.82%, correcting 91.7% of the numerical calculation errors. Furthermore, we isolate the knowledge reasoning capabilities of DeepSeek-R1 in complex financial reasoning tasks, which outperform other LLMs. More analyses are shown in Appendix F.
@@
Based on an analysis of the Hard subset, we observe that the PoT exhibits a markedly stronger performance than the CoT in multi-step and complex numerical reasoning tasks. Specifically, PoT leverages structured code generation to reduce token consumption. Under similar or lower token usage, PoT achieves greater accuracy (Table 3). Moreover, its performance is on par with certain LRMs that utilize test-time scaling strategies. For example, GPT-4o, when prompted with PoT, consumes only 54k tokens to solve the Hard subset, whereas CoT requires 173k tokens. Detailed statistics on token consumption are provided in LABEL:tab:tokens.
As shown in LABEL:fig:cost_effective, LLMs with PoT prompting method not only significantly reduce token overhead during inference, but can also approach or match the performance of LRMs. DeepSeek-R1 with CoT achieves the similar accuracy on the Hard subset as GPT-4o with PoT, but consumes much more tokens (742k vs. 54k). This “token-for-accuracy” test-time scaling strategy allows LRMs to maintain high correctness by repeatedly verifying outcomes from multiple perspectives. However, the associated inference cost is prohibitively high. In contrast, PoT achieves performance comparable to LRMs for complex financial calculations, while offering notably better cost-effectiveness.
@@ 5 Related Work
The emergence of reasoning models such as OpenAI o1 OpenAI (2024d) and DeepSeek-R1 DeepSeek-AI et al. (2025) has significantly improved the performance of LLMs in complex reasoning tasks in domains such as code Jain et al. (2025); Chen et al. (2021a), math Mao et al. (2024); Lightman et al. (2024), and science Lu et al. (2024); Yue et al. (2024); Wang et al. (2024). Among these large reasoning models OpenAI (2024d, c, 2025); DeepSeek-AI et al. (2025); Team (2024); Team et al. (2025); Gemini (2025), OpenAI o1 and DeepSeek-R1 have achieved competitive optimal performance. However, there currently exists a gap between the evaluation of LRMs and real-world domain-specific reasoning tasks, with a lack of evaluation and research on the model’s ability to flexibly apply domain knowledge in complex multi-step numerical reasoning. In the financial domain, the difficulty of questions and the quality of annotations become key limitations in evaluating the real reasoning capabilities of LRMs. For example, CodeFinQA and CodeTAT-QA Krumdick et al. (2024), which are derived from the classic financial question answering datasets FinQA Chen et al. (2021b) and TAT-QA Zhu et al. (2021), rely on tabular data extraction and simple arithmetic operations that cannot accurately assess the improvements in reasoning ability of LRMs compared to LLMs. For datasets such as FinCode Krumdick et al. (2024) and FinanceMath Zhao et al. (2024), limited complex problems, ambiguous questions, and relaxed evaluation criteria hinder the accurate assessment.
@@ 6 Conclusion
In this work, we introduce FinanceReasoning, a benchmark dataset designed to evaluate and enhance large language models in complex financial numerical reasoning tasks that require multi-step quantitative analysis, precise formula application, and hybrid contextual understanding. However, there are still some limitations: (1) We process tabular content as text, whereas in real-world scenarios, tables may also appear as images, requiring additional processing steps. In such cases, datasets such as MathVista Lu et al. (2024) and MMMU Yue et al. (2024), which focus on reasoning over image-based questions, serve as valuable complements to our benchmark. We believe that incorporating elements from these datasets into FinanceReasoning could help bridge the gap between text-based and multimodal financial reasoning, enabling a more comprehensive assessment of LLMs’ real-world applicability. (2) Due to limited resources, we do not conduct an evaluation of OpenAI o1 with PoT on the Easy and Medium subsets, as preliminary experiments suggest they are less challenging, and existing LLMs already demonstrate strong performance on these levels. (3) While we systematically verified and updated numerical answers and program solutions for multiple published datasets, we were unable to perform the same verification for the 1,000-problem test subset of FinanceMath Zhao et al. (2024), as it does not publicly provide ground-truth references or Python solutions, limiting our ability to ensure consistency in result validation. We commit to incorporating additional FinanceMath test subset should it become publicly available in the future, subject to the same rigorous validation process. (4) While FinanceReasoning provides comprehensive evaluation criteria, its long-term utility depends on adapting to evolving challenges. To address this, we commit to maintaining and periodically updating the benchmark through versioned releases on Hugging Face Datasets. (5) We clarify that FinanceReasoning evaluates perfect-information scenarios where all solution prerequisites are explicitly provided, thus focusing on the upper bounds of LRMs’ precision in deterministic calculations. This design intentionally distinguishes reasoning errors (e.g., outputting -3.1 versus 310,000 given clear constraints) from ambiguity resolution capabilities. While the important direction of modeling LRMs’ proactive clarification-seeking behavior under insufficient conditions falls outside this work’s scope, we will pursue it in future research through controlled ambiguity injection and interactive evaluation protocols.

### v2

@@ Abstract
Dataset Subset Answer Corrections Question Disambiguations Total
@@
Dataset (Criteria) DeepSeek-V3 DeepSeek-R1 $\Delta$ (R1 - V3) % Relative Increase (R1 vs. V3)
@@
Table 1: (a) Re-annotation: For the test or validation sets of four datasets, the proportion of updated examples ranges from 9.72% to 30.00%. (b) Re-evaluation: Silver denotes the Accuracy on the original dataset under relaxed criteria, while Gold represents the results on the re-annotated dataset under strict criteria. Rigorous revision and evaluation reveal LRMs’ actual performance and DeepSeek-R1’s significant improvement over DeepSeek-V3.
@@ 2.1 Updates to Public Datasets
Following prior work Krumdick et al. (2024); Zhao et al. (2024), we retain the format of questions with optional hybrid contexts as input, accompanied by Python-formatted solutions and program-executed numerical results. Due to the specialized nature of complex financial problems and the high cost of expert annotation, we observe certain limitations in existing datasets Chen et al. (2021b); Zhu et al. (2021); Zhao et al. (2024), including unsolvable problems, ambiguous statements (e.g., the phrasing “the range of” confuses the LLMs, as it is not clear whether to output a range like 70.18-81.05 or a specific difference of 10.87), oversimplified processes, incorrect answers, and relaxed evaluation criteria. Statistical details of these issues are presented in Table 1.
@@ 2.2 Function Library Construction
We begin by collecting 6,138 financial encyclopedia articles from Investopedia, a platform renowned for its extensive expertise in financial knowledge [^1] https://www.investopedia.com. Each article provides a detailed introduction to a specific financial term, covering fundamental concepts, application scenarios, and potential limitations, some including relevant calculation formulas and practical examples. To distill dense, structured financial reasoning knowledge while reducing annotation costs, we instruct GPT-4o to extract potential financial calculation functions from each article according to a predefined format. Each function is required to include a semantically meaningful signature, a concise and clear docstring (functionality, parameters, return values, applicable constraints, and other notes), and step-by-step implementation code with appropriate annotations. Finally, we organize financial experts to rigorously review and revise the generated functions, ensuring their professional expression and logical correctness. Detailed examples of financial functions are illustrated in Figure 3 and LABEL:tab:exmFunc.
@@ 2.3 Expansion of Data Annotation
Existing financial question-answering datasets (e.g., CodeFinQA, CodeTAT-QA Krumdick et al. (2024)) focus primarily on evaluating models’ basic concept understanding, precise numerical extraction, and simple calculation abilities within given contexts. Problem-solving processes in these datasets typically involve fewer reasoning steps (e.g., calculate the difference in net profit over two years). These datasets often suffer from redundancy in simple questions and a lack of complex questions, failing to adequately assess the reasoning capabilities of LRMs, such as knowledge application, constraint emphasis and long thought (e.g., compute the net return rate of a fund in Figure 1). As a result, the real reasoning capabilities of LRMs cannot be evaluated comprehensively and objectively. Therefore, optimizing data construction methods, rigorously verifying data quality, and building more challenging datasets have become crucial to improve the evaluation of financial reasoning tasks.
@@ 2.4 Data Quality Assurance
Table 2: Statistics of the financial function library and FinanceReasoning dataset (Avg values of three subsets).
@@ 2.5 Data Grading and Statistics
As illustrated in Table 2 and Figure 4, based on the difficulty of reasoning, we divided the problems into three subsets: Easy (1,000 examples), Medium (1,000 examples), and Hard (238 examples). More analyses are in Appendix C.2. To promote the evaluation of LRMs’ reasoning capabilities in the financial domain, we have made all problems publicly available, along with the complete financial function library.
@@ 3 Evaluation System
We developed an evaluation system for complex reasoning on FinanceReasoning, where all evaluations of LLMs were conducted by calling their official API interfaces. LABEL:tab:model-detail illustrates the exact model versions we used.
@@ 3.1 Large Language Models
We focused on evaluating six of the most powerful LRMs currently available.
@@ 3.2 Evaluation Methods
Following Zhao et al. (2024), we evaluated LLMs with CoT Wei et al. (2022) and PoT Chen et al. (2023b) to achieve optimal performance and make comparisons.
@@
We adopt the answer extraction pipeline from Zhao et al. (2024), using GPT-4o-mini to extract numerical results from the output in CoT setting, and executing the program from the output in PoT setting. Finally, we perform a strict accuracy evaluation comparing the numerical results with the ground truth within 0.2% error margin.
@@ 4.1 Main Results (RQ1)
The results demonstrate that the powerful LRM (i.e., OpenAI o1) using PoT prompting method achieves the best performance, with an accuracy of 89.1% on the Hard subset, significantly outperforming other LRMs and LLMs. On the Easy and Medium subsets, the evaluated LLMs achieve accuracy above 87%, except for the Llama models, where the advantage of LRMs is less pronounced. This further validates that simpler datasets have largely been solved by LLMs, making it difficult to assess the real reasoning capabilities of LRMs. On the Hard subset, LRMs exhibit a clear advantage over LLMs with CoT, further confirming the superiority of LRMs in complex reasoning tasks.
In particular, on the Hard subset, the current superior LRMs (i.e., OpenAI o1 and DeepSeek-R1) exhibit distinct performance contrasts in CoT and PoT settings. In CoT, DeepSeek-R1 achieves a 2.1% higher accuracy than OpenAI o1 (83.2% vs. 81.1%). However, the PoT prompting method significantly enhances OpenAI o1’s performance, allowing it to surpass DeepSeek-R1 (89.1% vs. 85.3%). This suggests that DeepSeek-R1 excels at text-based step-by-step reasoning, while OpenAI o1 outperforms in programming capabilities. This discrepancy may be due to differences in their training methods and training data.
@@ 4.2 Error Analysis (RQ2)
To better analyze the capabilities and limitations of LRMs on difficult problems in our dataset, we conduct a thorough and comprehensive error analysis. This analysis is based on 80 DeepSeek-R1 failure cases with PoT, with stratified sampling (20 Easy /20 Medium /40 Hard). We summarize four types of error in the current LRMs on challenging domain-specific reasoning problems, some of which involve compound errors. The detailed error distribution is shown in Table 4. More details of error cases are provided in Appendix B.
@@
  • Formula Application Errors and Numerical Calculation Errors increase with difficulty, especially the former, which supports the effectiveness of our difficulty categorization. In contrast, Numerical Extraction Errors remain consistently low across all subsets. These findings suggest that the model’s domain-specific knowledge comprehension and application capabilities should be enhanced through retrieval-augmented generation (RAG) or reinforcement learning (RL), thereby improving its performance on complex reasoning tasks in specialized domains.
@@ 4.3 Knowledge Augmentation (RQ3)
LLM-instructed Retrieval &
LLM as Judge
91.60 (+8.41) 86.97 (+3.78)
Table 5: Results of different knowledge augmentation methods on the Hard subset of FinanceReasoning. GPT-4o with refined knowledge augmentation, outperforming OpenAI o1 (91.6% vs. 89.1%) in PoT setting.
@@
As shown in Table 5, the format and method of knowledge augmentation significantly affect the performance of the model reasoning. Specifically, LLMs enhanced with financial function knowledge significantly outperform those enhanced with passage knowledge, as financial functions serve as refined reasoning knowledge. Excessive and intricate passages can disrupt the model’s reasoning abilities, resulting in diminished performance for both LLMs and LRMs. Taking advantage of the improved retrieval efficiency caused by LLM-Instructed Knowledge Retrieval, the combination approach achieves the best performance, improving the accuracy of GPT-4o to 91.6% with PoT. More analyses are shown in Appendix E.
@@ 4.4 Reasoner with Programmer (RQ4)
Specifically, we employ DeepSeek-R1 as Reasoner, and the best-performing models in PoT setting, Claude 3.5 Sonnet and GPT-4o, as Programmers, respectively. As shown in Figure 5, the combination of models achieves significant improvements compared to individual models. Compared to GPT-4o, Claude 3.5 Sonnet demonstrates a stronger ability to follow the given reasoning logic and generate precise code without introducing noise, owing to its programming advantages. The combination of DeepSeek-R1 and Claude 3.5 Sonnet achieves an accuracy of 87.82%, correcting 91.7% of the numerical calculation errors. Furthermore, we isolate the knowledge reasoning capabilities of DeepSeek-R1 in complex financial reasoning tasks, which outperform other LLMs. More analyses are shown in Appendix F.
@@
Based on an analysis of the Hard subset, we observe that PoT exhibits a markedly stronger performance than CoT in multi-step and complex numerical reasoning tasks. Specifically, PoT leverages structured code generation to reduce token consumption. Under similar or lower token usage, PoT achieves greater accuracy (Table 3). Moreover, its performance is on par with certain LRMs that utilize test-time scaling strategies. For example, GPT-4o, when prompted with PoT, consumes only 54k tokens to solve the Hard subset, whereas CoT requires 173k tokens. Detailed statistics on token consumption are provided in LABEL:tab:tokens.
As shown in LABEL:fig:cost_effective, LLMs using PoT prompting method not only significantly reduce token overhead during inference, but can also approach or match the performance of LRMs. DeepSeek-R1 with CoT achieves the similar accuracy on the Hard subset as GPT-4o with PoT, but consumes much more tokens (742k vs. 54k). This “token-for-accuracy” test-time scaling strategy allows LRMs to maintain high correctness by repeatedly verifying outcomes from multiple perspectives. However, the associated inference cost is prohibitively high. In contrast, PoT achieves performance comparable to LRMs for complex financial calculations, while offering notably better cost-effectiveness.
@@ 5 Related Work
The emergence of reasoning models such as OpenAI o1 OpenAI (2024d) and DeepSeek-R1 DeepSeek-AI et al. (2025) has significantly improved the performance of LLMs in complex reasoning tasks in domains such as code Jain et al. (2025); Chen et al. (2021a), math Mao et al. (2024); Lightman et al. (2024), and science Lu et al. (2024); Yue et al. (2024); Wang et al. (2024). Among these LRMs OpenAI (2024d, c, 2025); DeepSeek-AI et al. (2025); Team (2024); Team et al. (2025); Gemini (2025), OpenAI o1 and DeepSeek-R1 have achieved competitive optimal performance. However, there currently exists a gap between the evaluation of LRMs and real-world domain-specific reasoning tasks, with a lack of evaluation and research on the model’s ability to flexibly apply domain knowledge in complex multi-step numerical reasoning. In the financial domain, the difficulty of questions and the quality of annotations become key limitations in evaluating the real reasoning capabilities of LRMs. For example, CodeFinQA and CodeTAT-QA Krumdick et al. (2024), which are derived from the classic financial question-answering datasets FinQA Chen et al. (2021b) and TAT-QA Zhu et al. (2021), rely on tabular data extraction and simple arithmetic operations that cannot accurately assess the improvements in reasoning ability of LRMs compared to LLMs. For datasets such as FinCode Krumdick et al. (2024) and FinanceMath Zhao et al. (2024), limited complex problems, ambiguous questions, and relaxed evaluation criteria hinder the accurate assessment.
@@ 6 Conclusion
In this work, we introduce FinanceReasoning, a benchmark dataset designed to evaluate and enhance LLMs in complex financial numerical reasoning tasks that require multi-step quantitative analysis, precise formula application, and hybrid contextual understanding. However, there are still some limitations: (1) We process tabular content as text, whereas in real-world scenarios, tables may also appear as images, requiring additional processing steps. In such cases, datasets such as MathVista Lu et al. (2024) and MMMU Yue et al. (2024), which focus on reasoning over image-based questions, serve as valuable complements to our benchmark. We believe that incorporating elements from these datasets into FinanceReasoning could help bridge the gap between text-based and multimodal financial reasoning, enabling a more comprehensive assessment of LLMs’ real-world applicability. (2) Due to limited resources, we do not conduct an evaluation of OpenAI o1 with PoT on the Easy and Medium subsets, as preliminary experiments suggest they are less challenging, and existing LLMs already demonstrate strong performance on these levels. (3) While we systematically verified and updated numerical answers and program solutions for multiple published datasets, we were unable to perform the same verification for the 1,000-problem test subset of FinanceMath Zhao et al. (2024), as it does not publicly provide ground-truth references or Python solutions, limiting our ability to ensure consistency in result validation. We commit to incorporating additional FinanceMath test subset should it become publicly available in the future, subject to the same rigorous validation process. (4) While FinanceReasoning provides comprehensive evaluation criteria, its long-term utility depends on adapting to evolving challenges. To address this, we commit to maintaining and periodically updating the benchmark through versioned releases on Hugging Face Datasets. (5) We clarify that FinanceReasoning evaluates perfect-information scenarios where all solution prerequisites are explicitly provided, thus focusing on the upper bounds of LRMs’ precision in deterministic calculations. This design intentionally distinguishes reasoning errors (e.g., outputting -3.1 vs. 310,000 given clear constraints) from ambiguity resolution capabilities. While the important direction of modeling LRMs’ proactive clarification-seeking behavior under insufficient conditions falls outside this work’s scope, we will pursue it in future research through controlled ambiguity injection and interactive evaluation protocols.